Tohban Report 2010-11-03
From RHESSI Wiki
|Start Date:||27 Oct 2010|
|End Date:||03 Nov 2010|
|Next Tohban:||Juan Carlos Martinez Oliveros|
|List all reports|
There have been four ARs on disk throughout the week. However, NOAA 11120 and NOAA 11117 have been the most active, producing the vast majority of events. The former NOAA 11112 is approaching the east limb. Having produced an M2.9 flare, along with 7 C class flares during its last passage, B and C class events are probable from this region in the coming week, along with continued C class activity from NOAA 11120. Some of this weeks B class flares contain a significant HXR component e.g. on Oct 31st.
How many GOES flares occurred? Flares above B, C, M, X class were 83 4 0 0 And how many of these are listed in the RHESSI flare list? Flares above B, C, M, X class were 71 3 0 0 And how many had EXCELLENT coverage? Flares above B, C, M, X class were 71 2 0 0
The SSR is emptying out at 0% and is reaching a maximum at ~30%.
There were three VC3 data gaps:
GAPS IN APP_ID = 102 (VC3-MONITOR RATES) WITH PACKET RATE LT 30 N_GAPS 1 GAP START TIME GAP END TIME GAP (SEC) 2010-10-28T18:25:00.000 -- 2010-10-28T18:30:00.000 300.00000 GAPS IN APP_ID = 102 (VC3-MONITOR RATES) WITH PACKET RATE LT 30 N_GAPS 1 GAP START TIME GAP END TIME GAP (SEC) 2010-10-29T02:35:00.000 -- 2010-10-29T02:40:00.000 300.00000 GAPS IN APP_ID = 102 (VC3-MONITOR RATES) WITH PACKET RATE LT 30 N_GAPS 1 GAP START TIME GAP END TIME GAP (SEC) 2010-11-02T23:55:00.000 -- 2010-11-03T00:00:00.000 300.00000
No commanding changes this week. Spin rate is 14.9 and we remain in normal/vigorous decimation mode, with 4 minutes of data taken on each side of the eclipse. The PTP (programmable Telemetry Processor) has been replaced as it was causing problems. This is the original that has been used since launch. The replacement now takes two copies of both the high time res and SMEX headers - the primary copy and the backup.